The temperature in space given by T(x, y, z)= 200xyz? Find the hottest temperature on a unit sphere centered at the origin. (DO NOT USE LAGRANGE MULTIPLIERS)

1
Expert's answer
2021-12-16T09:56:23-0500

unit sphere centered at origin is "x^{2}+y^{2}+z^{2}=1"

 "\Rightarrow z^{2}=1-x^{2}-y^{2}"

given "T(x, y, z)=200 x y z^{2}"

"T(x, y)=200 x y\left(1-x^{2}-y^{2}\right)"

"\Rightarrow T(x, y)=200\left(x y-x^{3} y-x y^{3}\right)"

"T_{x}=200\left(y-3 x^{2} y-y^{3}\right), T_{y}=200\left(x-x^{3}-3 x y^{2}\right)"

for critical points "T_{x}=0, T_{y}=0"

 "\begin{aligned}

&200\left(y-3 x^{2} y-y^{3}\right)=0,200\left(x-x^{3}-3 x y^{2}\right)=0 \\

&\Rightarrow 200 y\left(1-3 x^{2}-y^{2}\right)=0,200 x\left(1-x^{2}-3 y^{2}\right)=0 \\

&\Rightarrow y=0,1-3 x^{2}-y^{2}=0 ; x=0,1-x^{2}-3 y^{2}=0 \\

&y=0,1-x^{2}-3 y^{2}=0 \\

&\Rightarrow 1-x^{2}-3(0)^{2}=0 \\

&\Rightarrow x=-1,1

\end{aligned}"

 "\begin{aligned}

&x=0,1-3 x^{2}-y^{2}=0 \\

&\Rightarrow 1-3\left(0^{2}\right)-y^{2}=0 \\

&\Rightarrow y=-1,1 \\

&1-3 x^{2}-y^{2}=0,1-x^{2}-3 y^{2}=0 \\

&\Rightarrow 3-9 x^{2}-3 y^{2}=0,1-x^{2}-3 y^{2}=0 \\

&\Rightarrow 3-9 x^{2}-3 y^{2}-1+x^{2}+3 y^{2}=0-0 \\

&\Rightarrow 2-8 x^{2}=0 \\

&\Rightarrow x^{2}=1 / 4 \\

&\Rightarrow x=-1 / 2,1 / 2

\end{aligned}"

 "\begin{aligned}
&1-3 x^{2}-y^{2}=0, x^{2}=1 / 4 \\
&\Rightarrow 1-3(1 / 4)-y^{2}=0 \\
&\Rightarrow y^{2}=1 / 4 \\
&=y=-1 / 2,1 / 2 \\
&(-1,0),(1,0),(0,-1),(0,1),(-1 / 2,-1 / 2),(1 / 2,1 / 2),(-1 / 2,1 / 2),(1 / 2,-1 / 2)\\& \text { are only critical points on unit sphere } \\
&T(-1,0)=0, \\
&T(1,0)=0, \\
&T(0,-1)=0, \\
&T(0,1)=0, \\
&T(-1 / 2,-1 / 2)=200(-1 / 2)(-1 / 2)\left(1-(-1 / 2)^{2}-(-1 / 2)^{2}\right)=25, \\
&T(1 / 2,1 / 2)=200(1 / 2)(1 / 2)\left(1-(1 / 2)^{2}-(1 / 2)^{2}\right)=25 \\
&T(-1 / 2,1 / 2)=200(-1 / 2)(1 / 2)\left(1-(-1 / 2)^{2}-(1 / 2)^{2}\right)=-25, \\
&T(1 / 2,-1 / 2)=200(1 / 2)(-1 / 2)\left(1-(1 / 2)^{2}-(-1 / 2)^{2}\right)=-25 \\
&\text { hottest temperature on unit sphere is } 25 \text { units }
\end{aligned}"